OpenStudy (anonymous):

7(n+8)=98 show steps please. answer:6

OpenStudy (anonymous):

7(n+8)=98 Divide each term in the equation by 7. n+8=14 Move all terms not containing n to the right-hand side of the equation. n=6

OpenStudy (anonymous):

\[\Huge\infty \]

OpenStudy (anonymous):

7(n+8)=98 7n+56=98 7n=98-56 7n=42 n=42/7 n=6